How do I find the angle's measures in a scalene triangle with out a protractor? The only information given is that in the triangle, the 3 angles are: x+16, x+21 and x+35. So, I need to find the value of x using an algebraic equation. So, I have tried this: 180=(x+16)+(x+35)+(x=21) Because all angles in a triangle must add up to 180 degrees. But, the problem is, I dont know how to solve this equation and what to do. I've been told to isolate the variable. Im in grade 8. Please help, Thanks
Answer by richard1234(7193)

You're at a good start. Now it's simply an algebra problem. Collect the x terms, then collect the constant terms to obtain

180 = (x+16) + (x+35) + (x+21)

180 = 3x + 72. Subtract 72 from both sides:

108 = 3x, divide both sides by 3 to get x = 36. However, finding the value of x alone is not sufficient since we want the angle measures, so we substitute x into our angles:

x+16 = 36 + 16 = 52 (degrees)
x+21 = 57
x+35 = 71